OpenStudy (anonymous):

Write a possible explicit rule for the nth term of the sequence: 25, 15, 5, -5, -15,

OpenStudy (anonymous):

Seems to start at 25 and decrease by 10 each step.\[a_n = 25-10n\]

OpenStudy (anonymous):

but the 1st term is 25. So if you plug in 1 for n, it doesn't work. I'd say 25 - 10(n-1)

OpenStudy (anonymous):

if you simplify this it is 35-10n

OpenStudy (anonymous):

well my options are A) an = 35 - 10n B) an = 35(an-1) - 10 C)an = 35n - 10 D) an = 35 - 10(an-1)

OpenStudy (anonymous):

yeah i was going w/ A

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Id say thats correct !!

OpenStudy (anonymous):

yw